OpenStudy (anonymous):

use product-to-sum to write as a sum or difference 12sin(pi/8)cos(pi/8)

OpenStudy (anonymous):

\[\sin(2\alpha)=2\sin(\alpha)\cos(\alpha)\] here \(\alpha =\frac{\pi}{8}\) so \(2\alpha =\frac{\pi}{4}\)

OpenStudy (anonymous):

so you have \[6\times 2\sin(\frac{\pi}{8})\cos(\frac{\pi}{8})\] \[=6\times \sin(\frac{\pi}{4})=6\times \frac{\sqrt{2}}{2}=3\sqrt{2}\]

OpenStudy (anonymous):

Okay, and you're 100% sure of this ? you how did you get the radical 2/2 ? from the unit circle ?

OpenStudy (anonymous):

because i recall what \(\sin(\frac{\pi}{4})\) is as it is a fairly common one. but you can check the unit circle on the last page of the cheat sheet i am attaching and you will see that \(\sin(\frac{\pi}{4})=\frac{\sqrt{2}}{2}\)

OpenStudy (anonymous):

you seem skeptical (which is a good thing) did i make a mistake?

OpenStudy (anonymous):

nope, it seems good(: I just wanna make sure I understand it correctly. Nothing worst than learning something wrong(:

OpenStudy (anonymous):

Thank you. I'm posting 2 more like this and a verifying identity if you wanna take a look in a minute(:

OpenStudy (anonymous):

you know what they say, "trust, yet verify" http://www.wolframalpha.com/input/?i=12sin%28pi%2F8%29cos%28pi%2F8%29

OpenStudy (anonymous):

haha true, thank you :D